1. Suppose that the demand for crude oil is given by Qd =85- 0.4P, where Qd is the quantity demanded in millions of barrels per day and P is the price per barrel in dollars. Suppose also that the supply of crude oil is given by Qs= 55 + 0.6P.

a) Calculate equilibrium price and quantity in this Market

b) Graph the demand curve and the supply curve

c) Calculate the demand for crude oil and the supply for crude oil if the market price is $ 15.00 per barrel. Explain your answer.

d) Calculate the demand for crude oil and the supply for crude oil if the market price is $50.00 per barrel. Explain your answer.

2. What type of relationship exists between the firm’s average physical product curve and the average cost curve in the short- run?


1
Expert's answer
2021-10-05T09:20:39-0400

Solution:

a.). At equilibrium: Qd = Qs

85 – 0.4P = 55 + 0.6P

85 – 55 = 0.6P + 0.4P

30 = P

Equilibrium Price = 30

Substitute in either the demand or supply function to derive Quantity:

85 – 0.4(30) = 85 – 12 = 73

Equilibrium Quantity = 73 units

 

b.). The demand and supply curve for crude oil is as below:

 


 

c.). If the market price is $15, the demand and supply for crude oil will be as follows:

Qd = 85 – 0.4(15) = 85 – 6 = 79 units

Qs = 55 + 0.6(15) = 55 + 9 = 64 units

This is because the market for crude oil is price takers and hence will adopt the market price rather than the equilibrium price. Since the price will be lower than the equilibrium price, the quantity demanded will increase. Producers, on the other hand, will supply less quantity due to the reduction in oil prices.

 

d.). Qd = 85 – 0.4(50) = 85 – 20 = 65 units

Qs = 55 + 0.6(50) = 55 + 30 = 85 units

Since the price is above the equilibrium price, the quantity demanded will decrease while the quantity supplied will increase since the higher prices stimulate producers to produce more.

 

2.). At the point of intersection between APP and AVC in the short-run, the firm would be producing at its shutdown point.